OpenStudy (anonymous):

Help with negative exponents please: (c^-1 + d^-1) \ (c^-2 + d^-2)

jimthompson5910 (jim_thompson5910):

hint: multiply every term by c^2d^2

OpenStudy (anonymous):

I did that and got cd (d + c) \ (d^2 + c^2)...can this be simplified?

jimthompson5910 (jim_thompson5910):

I'm getting \[\large \frac{cd(d+c)}{d^2 + c^2}\] which is what you have and you can't simplify further so nice work

OpenStudy (anonymous):

thank you so much! are you available to help with another problem?

jimthompson5910 (jim_thompson5910):

sure I can do one more

OpenStudy (anonymous):

thank you :) (x^-3 + y^-3) \ (x^-1 + y^-1)

jimthompson5910 (jim_thompson5910):

multiply everything by x^3y^3 and tell me what you get

OpenStudy (anonymous):

(y^3 + x^3) \ x^2y^3

jimthompson5910 (jim_thompson5910):

multiply each term by x^3y^3 to get x^-3 times x^3y^3 = y^3 y^-3 times x^3y^3 = x^3 so that part is correct in the numerator

jimthompson5910 (jim_thompson5910):

in the denominator we have x^-1 times x^3y^3 = x^2y^3 y^-1 times x^3y^3 = x^3y^2

jimthompson5910 (jim_thompson5910):

so the final answer is \[\large \frac{y^3+x^3}{x^2y^3+x^3y^2}\] and optionally you can factor to get \[\large \frac{y^3+x^3}{x^2y^2(y+x)}\]

OpenStudy (gabylovesyou):

hi @jim_thompson5910 did u get my PM?

jimthompson5910 (jim_thompson5910):

oh I guess you can factor the numerator as well to get y^3 + x^3 = (y+x)(y^2 + yx + x^2)

OpenStudy (anonymous):

how do you get the + x^3y^2 in the denominator?

OpenStudy (anonymous):

oh i see that in the numerator... would that be the same as (x + y) (x^2 + xy + y^2)?

jimthompson5910 (jim_thompson5910):

in the denominator, we have the terms x^-1 + y^-1 multiply those term by x^3y^3 to get x^2y^3+x^3y^2

OpenStudy (anonymous):

sorry I see

jimthompson5910 (jim_thompson5910):

sorry made a typo, it should be y^3 + x^3 = (y+x)(y^2 - yx + x^2)

jimthompson5910 (jim_thompson5910):

so the full step by step picture looks like this

jimthompson5910 (jim_thompson5910):

\[\large \frac{x^{-3} + y^{-3}}{x^{-1} + y^{-1}}\] \[\large \frac{x^3y^3*x^{-3} + x^3y^3*y^{-3}}{x^3y^3*x^{-1} + x^3y^3*y^{-1}}\] \[\large \frac{y^3+x^3}{x^2y^3+x^3y^2}\] \[\large \frac{(y+x)(y^2 - yx + x^2)}{x^2y^2(y+x)}\] \[\large \frac{\cancel{(y+x)}(y^2 - yx + x^2)}{x^2y^2\cancel{(y+x)}}\] \[\large \frac{y^2 - yx + x^2}{x^2y^2}\] ------------------------------------------------------------------------------------------ Final Answer: \[\large \frac{y^2 - yx + x^2}{x^2y^2}\]

OpenStudy (anonymous):

thank you so much for taking the time to explain it so well!

jimthompson5910 (jim_thompson5910):

yw